OpenStudy (anonymous):

how to solve this equation? 4x^(1/3)=2x^(-2/3)

OpenStudy (anonymous):

could you please use Equation

OpenStudy (anonymous):

otherwise it looks messy ` thanks

OpenStudy (ghazi):

\[4x^{1/3}=2x^{-2/3}\] it becomes \[2X^{\frac{ 1 }{ 3 }+\frac{ 2 }{ 3 }}=2X^\frac{ 3 }{ 3 }=2x\]

OpenStudy (anonymous):

i dont get it ?

OpenStudy (ghazi):

hold on

OpenStudy (ghazi):

well that 2x=1 and x= 1/2 , just keep the x and its operation on

OpenStudy (anonymous):

got it `

OpenStudy (anonymous):

thanks `

OpenStudy (ghazi):

\[4x^\frac{ 1 }{ 3 }=2x^\frac{ -2 }{ 3 }\] now \[\frac{ 4 x^\frac{ 1 }{ 3 }}{ 2x^\frac{ -2 }{ 3 } }=2x^{\frac{ 1 }{ 3 }+\frac{ 2 }{ 3 }}\]

OpenStudy (ghazi):

:D YW